A ball is thrown vertically upwards.When it is half of its maximum height its velocity is 10 metre per second ,find the maximum height reached by ball.

Grade:9

2 Answers

Vikas TU
4 years ago
Hiiiย 
V^2 = u ^2 + 2as
v = 0ย 
so , u^2 = 20Hย 
Now, apply this in middle , V = 10m/sย 
and s = H/2ย 
so height comes out to be 10mย 
ย 
sree lakshmi
4 years ago
Let us suppose that the ball was thrown with an initial speed u
Now, let us assume the Maximum Height reached is H
A constant acceleration g=10m/s2ย  is acting in downward direction.ย 

Let us consider the motion from start to Max Height.
In this time interval, final velocity is zero, and distance travelled is H. We have:
v^{2} = u^{2}+2as
\rightarrow 0^{2}=u^{2}+2\times (-10)H
u^{2}=20H
ย 
Now, let us consider the time from start to when the ball reaches half the maximum height. Now, in this time interval, final velocity is 10 m/s, and Distance Travelled is H/2
v^{2} = u^{2}+2as
10^{2}=20H +2(-10)H/2
100=20H-10H
10H=100
H= 10m
ย 
Thus, maximum height reached is 10 metres.